OpenStudy (anonymous):

laplace transform of f(t)=t^8e^0.5t

OpenStudy (anonymous):

try it and i will help you

OpenStudy (anonymous):

\[f(t) = t^8 * e ^{0.5t}\]

OpenStudy (anonymous):

\[L(f(t))=L(t^ne^{at})=\frac{n!}{(s-a)^{n+1}}\] that should fit your situation nicely. Unless you need to learn how to derive it from the integral form.

OpenStudy (anonymous):

Why isn't it 8!/s^9 * 1/(s-1/2) ? (I'm just reading it off the tables)